# Multiply Robust Estimation for Local Distribution Shifts with Multiple Domains

### v1

@@ Abstract
Distribution shifts are ubiquitous in real-world machine learning applications, posing a challenge to the generalization of models trained on one data distribution to another. We focus on scenarios where data distributions vary across multiple segments of the entire population and only make local assumptions about the differences between training and test (deployment) distributions within each segment. We propose a two-stage multiply robust estimation method to improve model performance on each individual segment for tabular data analysis. The method involves fitting a linear combination of the based models, learned using clusters of training data from multiple segments, followed by a refinement step for each segment. Our method is designed to be implemented with commonly used off-the-shelf machine learning models. We establish theoretical guarantees on the generalization bound of the method on the test risk. With extensive experiments on synthetic and real datasets, we demonstrate that the proposed method substantially improves over existing alternatives in prediction accuracy and robustness on both regression and classification tasks. We also assess its effectiveness on a user city prediction dataset from a large technology company.
@@ 1 Introduction
While building domain [^1] We use “domain” and “segment” interchangeably hereafter.-specific models for all segments and then adjusting for sampling bias may be a straightforward solution, it presents at least three challenges: (1) unreliable model training on segments with an insufficient number of labels, (2) excessive computation when training individual models with a large number of segments, and, (3) a potential loss of information shared across different segments. To address these challenges, we aim to propose a flexible modeling framework for general supervised learning scenarios on tabular data, with the goal of improving model performance on each individual domain.
@@
To make accurate predictions of $Y^{\prime}$ under sampling bias, it is necessary to assume that the distribution of the training data $P$ and that of the test data $Q$ share a persistent property, a component that is identical for $P$ and $Q$, to allow the model learned from the training data to be generalized to the test data. In classic domain adaptation literature, the two fundamental persistent properties are covariate shift (Shimodaira, 2000) and label shift (Saerens et al., 2002). In this work, we adopt a local assumption that such persistent property only holds within each segment, without positing any restrictive global assumptions on how data distributions are different across segments. We refer to this assumption as local distribution shift. The problem setup we have described above is visualized in Figure 1.
@@ 3 Problem Formulation
Unsupervised domain adaptation. In the setting of domain adaptation, we consider $N$ segments that span the whole population. Let $S\in\mathcal{S}=\{1,2,...,N_{S}\}$ denote the segment index, $\{Y_{i},X_{i},S_{i}\}_{i=1}^{n_{tr}}$ and $\{Y^{\prime}_{j},X^{\prime}_{j},S^{\prime}_{j}\}_{j=1}^{n_{te}}$ be the training and testing sets sampled independently from population probability distributions $P$ and $Q$ respectively, where $P\not=Q$. Our goal is to construct segment-specific predictors $f^{(s)}:\mathcal{X}\to\mathcal{Y}$ which perform well on each segment’s test distribution $Q^{(s)}\equiv Q(y,x|S=s)$ according to the segment-conditional risk $R_{Q}[f|s]=\mathbb{E}_{(X,Y)\sim Q^{(s)}}[\ell(Y,f(X))]$, where $\ell(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a loss function. We consider unsupervised domain adaptation where the test labels $Y^{\prime}_{j}$s are unobserved. We similarly define $P^{(s)}\equiv P(y,x|S=s)$.
Local distribution shifts. Classical assumptions include covariate shift (Shimodaira, 2000) where $P(y|x)=Q(y|x),~{}P(x)\not=Q(x)$ and label shift (Saerens et al., 2002) where $P(x|y)=Q(x|y),~{}P(y)\not=Q(y)$. In our context, we adopt a relaxed version of these shifts where either covariate shift or label shift holds within each segment throughout this work.
@@
  • Local label shift: $P^{(s)}(x|y)=Q^{(s)}(x|y)$
We highlight the motivation for these relaxed assumptions with a toy example. Consider a scenario where the manifestation of symptoms $X$ (such as fever) is contingent upon the status of an influenza infection $Y$, and we wish to develop a classifier for this year’s flu optimized for both children ($S=0$) and adults ($S=1$) that exhibit differences in symptoms of infection $P(x|y,s=0)\not=P(x|y,s=1)$. Here, $P$ represents the training distribution derived from historical data, while $Q$ signifies the test distribution for the current flu season. Across various flu seasons, the prior propensity of contracting the flu may vary, expressed as $P(y|s)\neq Q(y|s)$. However, the distribution of symptoms given a flu status remains consistent, i.e., denoted by $P(x|y,s)=Q(x|y,s)$, thereby illustrating a local label shift.
@@ 4 Multiply Robust Estimation
In this section we introduce our proposed multiply robust estimation method. This approach allows for information sharing across segments and adaptive adjustments for each domain, resulting in enhanced model performance on each specific domain. An overview of the proposal is as follows. We first cluster the segments into several groups and train a base model on each cluster. This enables information sharing across segments with similar data distributions and particularly beneficial for segments with a limited number of samples to leverage strength from others. To adjust for local distribution shifts, we estimate the importance weights to match the training data distribution to test data for each segment. We then learn a domain-specific model by finding the best linear combination of the base models, which we refer as stage 1 estimators. With the stage 1 estimators as “priors”, we further refine the estimators using weighted training samples. The pseudo-code of this procedure is presented in Algorithm 1
@@
  Step 2: Cluster the segments $\mathcal{C}_{M}(\{Y_{i},X_{i},S_{i}\}_{i=1}^{n_{tr}})=\{\mathcal{S}_{m}\}_{m=1}^{M}$ into $M-1$ clusters and a final group partition which includes all clusters ${\cal S}_{M}={\cal S}$
  Step 3: For each $m$ train a base model using the segments $h_{m}={\cal H}\left(\{Y_{i},X_{i}\}_{i\in N_{base}:S_{i}\in\mathcal{S}_{m}}\right)$
@@ 4.1 Two Stage Domain-adaptive Estimation
We now elaborate the two-stage estimation method. Let $\{h_{m}\}_{m=1}^{M},h_{m}:{\cal X}\to{\cal Y}$ denote a set of base models and $\mathbf{h}$ the vector collection of them. Let $R[f|Y,X,s;\alpha]$ denote the $\alpha$-weighted empirical training risk of segment $s$,
$R[f|Y,X,s;\alpha]=\frac{1}{n_{s}}\sum_{i\in N_{s}}\alpha(Y_{i},X_{i})\ell(Y_{i},f(X_{i}))$
where where $N_{s}$ denotes the indices of the data corresponding to segment $s$: $N_{s}\subset\{1,2,\dots n\},n_{s}=|N_{s}|$, and $Y,X$ denote the whole training data. We abbreviate the unweighted empirical risk $R[f|Y,X,s;\mathbf{1}]=R[f|Y,X,s]$. We let $\mathcal{B}\subseteq\mathbb{R}^{M}$ denote a convex subset for the parameters. Lastly, let $\Omega[f,g]$ denote a penalty function for the deviation of $f$ and $g$ (for example $\lVert f-g\rVert_{\mathcal{F}}$ for a function space ${\cal F}$, or $\lVert\beta_{f}-\beta_{g}\rVert_{2}$ for the parameters of a linear model), where $\Omega[f]=\Omega[f,0]$. Given a set of estimated weights on the test set $\widehat{w}_{s}$, and a regularization parameter $\lambda^{\prime}_{s}$ [^2] This can be any type of regularization, such as restricting the number of trees in an ensemble, or penalizing the coefficients in a linear model.. Steps 4b and 4c are presented as follows
$\widehat{\beta}^{(s)}$ $=\text{argmin}_{\beta\in{\cal B}}R[\beta^{\intercal}\mathbf{h}|Y,X,s]$
@@
$\widehat{f}^{(s)}_{MR}$ $=\text{argmin}_{f\in{\cal F}}R[f|Y,X,s;\widehat{w}_{s}]+\lambda_{s}^{\prime}\Omega[\widehat{\beta}^{(s)\intercal}\mathbf{h},f]$
@@ 4.1.1 Stage 1 Estimators
When ${\cal B}=\mathbb{R}^{M}$, the stage 1 estimator is a simple linear combination of base models. For example, if the loss function is the mean square loss this can be computed using linear regression or logistic regression when the $\ell$ is the cross entropy loss.
@@ 4.1.2 Stage 2 Estimators
The stage 2 estimator requires penalization towards a base model. In the case of the mean square loss and a form of the penalty $\Omega[f,g]=\Omega[f-g]$, then this can be computed by fitting a regularized model fit to the residuals of the stage 1 model, $\delta Y_{i}=Y_{i}-\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{i})$. For other loss functions, such as those involved in classification, one can include a base margin $\widehat{\delta}_{i}=\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{i})$ then one can learn a model $\tilde{g}$ such that $f=\tilde{g}+\widehat{\delta}_{i}$ by penalizing the complexity of $\tilde{g}$, through $\Omega[\tilde{g}]$. This functionality is available in popular packages such as XGBoost (Chen & Guestrin, 2016, XGB) and can be similarly implemented in many common function classes such as SVMs, tree-based methods and neural networks.
@@ 4.2.1 Segment Clustering for Base Model Training
We use Equation 1 to compute a distance matrix between segments and perform hierarchical clustering as described in Algorithm 2. The choice of kernel may be determined by the datatype. For example, we choose the delta kernel $k_{y}(y,y^{\prime})=1-I(y=y^{\prime})$ for categorical dimensions, where $I$ is the indicator function, and choose the Gaussian kernel for continuous dimensions. The joint kernel $k_{y,x}$ can be computed as the product of the $k_{x}$ and $k_{y}$ kernels, $k_{y,x}((y,x),(y^{\prime},x^{\prime}))=k_{y}(y,y^{\prime})k_{x}(x,x^{\prime})$.
@@ 4.3 Computational Complexity
The major computation burden in the proposed multiply robust method occurs at the segment clustering and 2-stage estimation. In Algorithm 2 for segment clustering, the distance matrix computation requires $\mathcal{O}(n_{tr}^{2})$ operations, which is typically the most expensive part. The agglomeration clustering step has complexity $\mathcal{O}(MN_{S}^{2})$ which tends to be much smaller as $N_{S}\ll n_{tr}$.
Suppose the complexities of training stage 1 and stage 2 estimators are $c_{1}(n_{s})$ and $c_{2}(n_{s})$ for segment $s$ in step 4 of Algorithm 1. The total algorithmic complexity is $\sum_{s=1}^{N_{S}}c_{1}(n_{s})+c_{2}(n_{s})$. For example, if we adopt linear regression in stage 1 and XGB in stage 2, then $c_{1}(n_{s})=\mathcal{O}(n_{s}M^{2}+M^{3})$ and $c_{2}(n_{s})=\mathcal{O}(TD_{T}dn_{s}+pn_{s}\log(n_{s}))$ resulting in the total complexity scaling log-linearly with $n_{tr}$. Here $T$ is the total number of trees and $D_{T}$ is the maximum tree depth and $d$ is the dimension of $X$. The base model training can be analyzed similarly as the stage 2 estimator, but for training $M$ models.
@@ 5 Theoretical Analysis
We now present a theoretical justification for the proposed multiply robust estimation method. As our goal is to construct a model that is adaptive to each segment with high predictive accuracy, we provide a performance guarantee by establishing a generalization bound on the test risk. Our proof leverages Theorem 1.1 from Ostrovskii & Bach (2021) to establish the convergence rate of the stage 1 model. We integrate a modified version of the risk minimization technique presented in Reddi et al. (2015). Notably, our methods are adaptable to any weight estimator, allowing us to separate the bound based on the quantities associated with the weight estimator $\widehat{w}_{s}$ and the function estimation of $\widehat{f}^{(s)}_{MR}$.
Recall the training risk for segment $s$, $R_{P}[f|s]=\mathbb{E}_{(Y,X)\sim P^{(s)}}[\ell(Y,f(X))]$ and denote the corresponding test risk $R_{Q}[f|s]=\mathbb{E}_{(Y,X)\sim Q^{(s)}}[\ell(Y,f(X))]$, and $\ell(y,\eta)$ is a loss function which is convex in $\eta$ for each $y$, such as the mean square loss for regression or cross entropy loss for classification. Our goal is to minimize the test risk $R_{Q}[f|s]$ for any given segment $s$. We focus on one segment to establish the theory, which can then be applied to all of the segments.
Suppose that the base models $\mathbf{h}=\{h_{m}\}_{m=1}^{M}$ are given and both base models and the candidate model $f$ are contained in a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{F}$ equipped with the norm $\lVert\cdot\rVert_{{\cal F}}$. Let $\beta^{*(s)}=\text{argmin}_{\beta\in{\cal B}}R_{P}[\beta^{\intercal}\mathbf{h}|s]$ be the optimal linear combination and denote the following quantities involving the base models
@@
Let $\ell^{\prime}(y,\eta)=\frac{d}{d\eta}\ell(y,\eta)$ and similarly for higher order derivatives. Assume that each $Q^{(s)}$ is absolutely continuous with respect to $P^{(s)}$ such that there exists a set of true importance weights $w_{s}(x,y)=\frac{dQ^{(s)}(x,y)}{dP^{(s)}(x,y)}$.
@@
$|\ell(y,\eta)|$ $\leq L,\quad|\ell(y,\eta)-\ell(y,\eta^{\prime})|\leq L|\eta-\eta^{\prime}|$ (2)
|′′′$|\ell^{\prime\prime\prime}(y,\eta)|$ |′′$\leq|\ell^{\prime\prime}(y,\eta)|,\quad\text{ for all }\eta\in\mathbb{R}$ (3)
@@
[…]
@@
[…]
$\frac{\lVert\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{s})\rVert_{2}}{n_{s}}\bigg{)}$
where $\lVert\widehat{w}_{s}-w_{s}\rVert_{1}=\sum_{i\in N_{s}}|\widehat{w}_{s}(X_{i},Y_{i})-w_{s}(X_{i},Y_{i})|$ and $K_{s}$ is the Gram matrix corresponding to the kernel function $k(\cdot,\cdot)$ associated with the space ${\cal F}$, $K_{s,ij}=k(X_{i},X_{j}):i,j\in N_{s}$, and $\lVert\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{s})\rVert_{2}=\sum_{i\in N_{s}}(\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{i}))^{2}$
@@
An alternative competitor one could suggest is to apply the DR estimator separately over every segment. Instead of employing a linear combination for the first stage, the DR learns an estimate $\widehat{f}^{(s)}_{p,\nu_{p,s}}\in{\cal F}$ trained on the unweighted empirical risk. Our comparison focuses on the functional complexity component of our bounds. They include a term analogous to our $\nu^{\prime}$ in their bound, which we denote as $\nu_{DR}^{\prime}$. However, they demonstrate $\nu_{DR}^{\prime}=\lVert f^{*(s)}_{\nu_{p,s}}-f^{*(s)}_{\nu_{q,s}}\rVert_{{\cal F}}+\mathcal{O}(n_{s}^{-1/4})$, where $f^{*(s)}_{\nu_{p,s}}$ represents the population optimizer of the unweighted problem. In contrast, our first stage estimator is expressed as $\nu^{\prime}=\lVert\beta^{*(s)\intercal}\mathbf{h}-f^{*(s)}_{\nu_{q,s}}\rVert_{\mathcal{F}}+\mathcal{O}(n_{s}^{-1/2})$, suggesting an improved rate for segments with small samples. Moreover if $\beta^{*(s)\intercal}\mathbf{h}$ proves to be a better approximation through similar segments to $f^{*(s)}_{\nu_{q,s}}$ then we can improve on the constant as well.

### v2

@@ Abstract
Distribution shifts are ubiquitous in real-world machine learning applications, posing a challenge to the generalization of models trained on one data distribution to another. We focus on scenarios where data distributions vary across multiple segments of the entire population and only make local assumptions about the differences between training and test (deployment) distributions within each segment. We propose a two-stage multiply robust estimation method to improve model performance on each individual segment for tabular data analysis. The method involves fitting a linear combination of the based models, learned using clusters of training data from multiple segments, followed by a refinement step for each segment. Our method is designed to be implemented with commonly used off-the-shelf machine learning models. We establish theoretical guarantees on the generalization bound of the method on the test risk. With extensive experiments on synthetic and real datasets, we demonstrate that the proposed method substantially improves over existing alternatives in prediction accuracy and robustness on both regression and classification tasks. We also assess its effectiveness on a user city prediction dataset from Meta.
@@ 1 Introduction
While building domain [^1] We use ‘domain’ and ‘segment’ interchangeably hereafter.-specific models for all segments and then adjusting for sampling bias may be a straightforward solution, it presents at least three challenges: (1) unreliable model training on segments with an insufficient number of labels, (2) excessive computation when training individual models with a large number of segments, and, (3) a potential loss of information shared across different segments. To address these challenges, we aim to propose a flexible modeling framework for general supervised learning scenarios on tabular data, with the goal of improving model performance on each individual domain; however, we also discuss how our methods could be adopted to image and text data.
@@
To make accurate predictions of $Y^{\prime}$ under distribution shift, it is necessary to assume that the distribution of the training data $P$ and that of the test data $Q$ share a persistent property, a component that is identical for $P$ and $Q$, to allow the model learned from the training data to be generalized to the test data. In classic domain adaptation literature, the two fundamental persistent property assumptions are covariate shift (Shimodaira, 2000) and label shift (Saerens et al., 2002). In this work, we adopt a local assumption that such persistent property only holds within each segment, without positing any restrictive global assumptions on how data distributions are different across segments. We refer to this assumption as local distribution shift. The problem setup we have described above is visualized in Figure 1.
@@ 3 Problem Formulation
Multi-Segment Unsupervised domain adaptation. In the setting of domain adaptation, we consider $N$ segments that span the whole population. Let $S\in\mathcal{S}=\{1,2,...,N_{S}\}$ denote the segment index, $\{Y_{i},X_{i},S_{i}\}_{i=1}^{n_{tr}}$ and $\{Y^{\prime}_{j},X^{\prime}_{j},S^{\prime}_{j}\}_{j=1}^{n_{te}}$ be the training and testing sets sampled independently from population probability distributions $P$ and $Q$ respectively, where $P\not=Q$. Our goal is to construct segment-specific predictors $f^{(s)}:\mathcal{X}\to\mathcal{Y}$ which perform well on each segment’s test distribution $Q^{(s)}\equiv Q(y,x|S=s)$ according to the segment-conditional risk $R_{Q}[f|s]=\mathbb{E}_{(X,Y)\sim Q^{(s)}}[\ell(Y,f(X))]$, where $\ell(\cdot,\cdot):\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a loss function. We consider unsupervised domain adaptation where the test labels $Y^{\prime}_{j}$s are unobserved. We similarly define $P^{(s)}\equiv P(y,x|S=s)$.
Local distribution shifts. Classical assumptions include covariate shift (Shimodaira, 2000) where $P(y|x)=Q(y|x),~{}P(x)\not=Q(x)$ and label shift (Saerens et al., 2002) where $P(x|y)=Q(x|y),~{}P(y)\not=Q(y)$. In our context, we adopt a relaxed version of these shifts where either covariate shift or label shift holds within each segment throughout this work. These assumptions include:
@@
  • Local label shift: $P^{(s)}(x|y)=Q^{(s)}(x|y)$.
We highlight the motivation for these relaxed assumptions with a toy example. Consider a scenario where the manifestation of symptoms $X$ (such as fever) is contingent upon the status of an influenza infection $Y$, and we wish to develop a classifier for this year’s flu optimized for both children ($S=0$) and adults ($S=1$) that exhibit differences in symptoms of infection $P(x|y,s=0)\not=P(x|y,s=1)$. Here, $P$ represents the training distribution derived from historical data, while $Q$ signifies the test distribution for the current flu season. Across various flu seasons, the prior propensity of contracting the flu may vary, expressed as $P(y|s)\neq Q(y|s)$. However, the distribution of symptoms given a flu status remains consistent, i.e., denoted by $P(x|y,s)=Q(x|y,s)$, thereby illustrating a local label shift as opposed to a global shift.
@@ 4 Multiply Robust Estimation
In this section we introduce our proposed multiply robust estimation method. This approach allows for information sharing across segments and adaptive adjustments for each domain, resulting in enhanced model performance on each specific domain. An overview of the proposal is as follows. We first cluster the segments into several groups and train a base model on each cluster. This enables information sharing across segments with similar data distributions and particularly beneficial for segments with a limited number of samples to leverage strength from others. To adjust for local distribution shifts, we estimate the importance weights to match the training data distribution to test data for each segment. We then learn a domain-specific model by finding the best linear combination of the base models, which we refer as stage 1 estimators. With the stage 1 estimators as ‘priors’, we further refine the estimators using weighted training samples. The pseudo-code of this procedure is presented in Algorithm 1.
@@
  Step 2: Cluster the segments $\mathcal{C}_{M}(\{Y_{i},X_{i},S_{i}\}_{i=1}^{n_{tr}})=\{\mathcal{S}_{m}\}_{m=1}^{M}$ into $M$ clusters and a final group partition which includes all clusters ${\cal S}_{M+1}={\cal S}$ where each $\mathcal{S}_{m}\subseteq\mathcal{S}$ are subsets of the set of segments.
  Step 3: For each $m$, train a base model using the segments $h_{m}={\cal H}\left(\{Y_{i},X_{i}\}_{i\in N_{base}:S_{i}\in\mathcal{S}_{m}}\right)$
@@
Remark 4.1.
In applications that involve image and text data, it is common to exploit large-scale pre-trained models such as VGG-16 (Simonyan & Zisserman, 2014). A straightforward and widely adopted approach to applying our method to text and image data is to utilize the embeddings of a large pre-trained model, treating them as the feature vector. Subsequently, our method can be applied, as outlined in Algorithm 1, employing a regularized, high-dimensional linear model for both the base-model and the stage 2 estimator classes. While addressing label shift issues typically involves straightforward application of BBSE, caution should be exercised when dealing with high-dimensional embeddings, necessitating careful selection of an appropriate covariate shift weight estimation method.
@@ 4.1 Two Stage Domain-adaptive Estimation
We now elaborate the two-stage estimation method. Let $\{h_{m}\}_{m=1}^{M},h_{m}:{\cal X}\to{\cal Y}$ denote a set of base models and $\mathbf{h}$ the vector collection of them. Let $R[f|\mathbf{Y},\mathbf{X},,s;\alpha]$ denote the $\alpha$-weighted empirical training risk of segment $s$,
$R[f|\mathbf{Y},\mathbf{X},s;\alpha]=\frac{1}{n_{s}}\sum_{i\in N_{s}}\alpha(Y_{i},X_{i})\ell(Y_{i},f(X_{i}))$
where $N_{s}$ denotes the indices of the data corresponding to segment $s$: $N_{s}\subset\{1,2,\dots n\},n_{s}=|N_{s}|$, and $\mathbf{Y},\mathbf{X}$ denote the entire training data. We abbreviate the unweighted empirical risk $R[f|\mathbf{Y},\mathbf{X},s;\mathbf{1}]=R[f|\mathbf{Y},\mathbf{X},s]$. We let $\mathcal{B}\subseteq\mathbb{R}^{M}$ denote a convex subset for the parameters. Lastly, let $\Omega[f,g]$ denote a penalty function for the deviation of $f$ and $g$ (for example $\lVert f-g\rVert_{\mathcal{F}}$ for a function space ${\cal F}$, or $\lVert\beta_{f}-\beta_{g}\rVert_{2}$ for the parameters of a linear model), where $\Omega[f]=\Omega[f,0]$. Given a set of estimated weights on the test set $\widehat{w}_{s}$, and a regularization parameter $\lambda^{\prime}_{s}$ [^2] This can be any type of regularization, such as restricting the number of trees in an ensemble, or penalizing the coefficients in a linear model.. Steps 4b and 4c are presented as follows
$\widehat{\beta}^{(s)}$ $=\text{argmin}_{\beta\in{\cal B}}R[\beta^{\intercal}\mathbf{h}|\mathbf{Y},\mathbf{X},s]$
@@
$\widehat{f}^{(s)}_{MR}$ $=\text{argmin}_{f\in{\cal F}}R[f|\mathbf{Y},\mathbf{X},s;\widehat{w}_{s}]+\lambda_{s}^{\prime}\Omega[\widehat{\beta}^{(s)\intercal}\mathbf{h},f]$
@@ 4.1.1 Stage 1 Estimators
When ${\cal B}=\mathbb{R}^{M}$, the stage 1 estimator is a simple linear combination of base models. For example, if the loss function is the mean square loss this can be computed using linear regression or logistic regression when the $\ell$ is the cross entropy loss. In practice, the performance of the model at this stage is influenced by the quality of the base models. Therefore, as with all machine learning problems, it is crucial to select a method that is suitable for the specific prediction task at hand, including kernel methods, ensemble methods, and others.
@@ 4.1.2 Stage 2 Estimators
The stage 2 estimator requires penalization towards a base model. In the case of the mean square loss and a form of the penalty $\Omega[f,g]=\Omega[f-g]$, then this can be computed by fitting a regularized model fit to the residuals of the stage 1 model, $\delta Y_{i}=Y_{i}-\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{i})$. For other loss functions, such as those involved in classification, one can include a base margin $\widehat{\delta}_{i}=\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{i})$ then one can learn a model $\tilde{g}$ such that $f=\tilde{g}+\widehat{\delta}_{i}$ by penalizing the complexity of $\tilde{g}$, through $\Omega[\tilde{g}]$. This functionality is available in popular packages such as XGBoost (Chen & Guestrin, 2016, XGB) and can be similarly implemented in many common function classes such as SVMs, tree-based methods and neural networks. In the case of XGBoost $\Omega[\tilde{g}]$ involves penalties on the complexity of the second stage model, i.e. the number of trees, tree depth etc. We illustrate a cross validation procedure for our method in Section C.1.
@@ 4.2.1 Segment Clustering for Base Model Training
We use Equation 1 to compute a distance matrix between segments and perform hierarchical clustering as described in Algorithm 2. The choice of kernel may be determined by the datatype. For example, we choose the delta kernel $k_{y}(y,y^{\prime})=1-I(y=y^{\prime})$ for categorical dimensions, where $I$ is the indicator function, and choose the Gaussian kernel for continuous dimensions. The joint kernel $k_{y,x}$ can be computed as the product of the $k_{x}$ and $k_{y}$ kernels, $k_{y,x}((y,x),(y^{\prime},x^{\prime}))=k_{y}(y,y^{\prime})k_{x}(x,x^{\prime})$. In our experiments, we generally adhere to a rule of thumb where we select the largest $M$ in Algorithm 2 that avoids generating singleton clusters, thereby ensuring a reasonably large function class in the base models. However, we note that generating an excessive number of distinct clusters could potentially restrict information sharing across segments, thereby diminishing the overall performance. One could also treat $M$ as a tuning parameter and select it through cross validation.
@@ 4.3 Computational Complexity
The major computation burden in the proposed multiply robust method occurs at the segment clustering and 2-stage estimation. In Algorithm 2 for segment clustering, the distance matrix computation requires $\mathcal{O}(n_{tr}^{2})$ operations, which is typically the most computationally expensive part of the procedure. The agglomeration clustering step has complexity $\mathcal{O}(MN_{S}^{2})$ which tends to be much smaller as $N_{S}\ll n_{tr}$.
Suppose the complexities of training stage 1 and stage 2 estimators are $c_{1}(n_{s})$ and $c_{2}(n_{s})$ for segment $s$ in step 4 of Algorithm 1. The total algorithmic complexity is $\sum_{s=1}^{N_{S}}c_{1}(n_{s})+c_{2}(n_{s})$. For example, if we adopt linear regression in stage 1 and XGB in stage 2, then $c_{1}(n_{s})=\mathcal{O}(n_{s}M^{2}+M^{3})$ and $c_{2}(n_{s})=\mathcal{O}(TD_{T}dn_{s}+pn_{s}\log(n_{s}))$ resulting in the total complexity scaling log-linearly with $n_{tr}$. Here $T$ is the total number of trees and $D_{T}$ is the maximum tree depth and $d$ is the dimension of $X$. The base model training can be analyzed similarly as the stage 2 estimator, but for training $M$ models. It is worth noting that each of the stage 1 and 2 estimators can be computed in parallel, also allowing for further algorithmic improvements in practice.
@@ 5 Theoretical Analysis
We now present a theoretical justification for the proposed multiply robust estimation method. As our goal is to construct a model that is adaptive to each segment with high predictive accuracy, we provide a performance guarantee by establishing a generalization bound on the test risk. Our proof leverages results from Ostrovskii & Bach (2021) to establish the convergence rate of the stage 1 model. We integrate a modified version of the risk minimization technique presented in Reddi et al. (2015). Notably, our methods are adaptable to any weight estimator, allowing us to separate the bound based on the quantities associated with the weight estimator $\widehat{w}_{s}$ and the function estimation of $\widehat{f}^{(s)}_{MR}$. The assumptions herein, though restrictive, are used to provide intuition on the performance of the algorithm and are not necessary for the algorithm to be effective in practice.
Recall the training risk for segment $s$, $R_{P}[f|s]=\mathbb{E}_{(Y,X)\sim P^{(s)}}[\ell(Y,f(X))]$ and denote the corresponding test risk $R_{Q}[f|s]=\mathbb{E}_{(Y,X)\sim Q^{(s)}}[\ell(Y,f(X))]$, and $\ell(y,\varrho)$ is a loss function which is convex in $\varrho$ for each $y$, such as the mean square loss for regression or cross entropy loss for classification. Our goal is to minimize the test risk $R_{Q}[f|s]$ for any given segment $s$. We focus on one segment to establish the theory, which can then be applied to all of the segments.
Suppose a set of base models $\mathbf{h}=\{h_{m}\}_{m=1}^{M}$ and the candidate model $f$ are contained in a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{F}$ equipped with the norm $\lVert\cdot\rVert_{{\cal F}}$. Let $\beta^{*(s)}=\text{argmin}_{\beta\in{\cal B}}R_{P}[\beta^{\intercal}\mathbf{h}|s]$ be the optimal linear combination and denote the following quantities derived from the base models
@@
Let $\ell^{\prime}(y,\varrho)=\frac{d}{d\varrho}\ell(y,\varrho)$ and similarly for higher order derivatives. Assume that each $Q^{(s)}$ is absolutely continuous with respect to $P^{(s)}$ such that there exists a set of bounded importance weights $w_{s}(x,y)=\frac{dQ^{(s)}(x,y)}{dP^{(s)}(x,y)}$.
@@
$|\ell(y,\varrho)|$ $\leq L,\quad|\ell(y,\varrho)-\ell(y,\varrho^{\prime})|\leq L|\varrho-\varrho^{\prime}|$ (2)
|′′′$|\ell^{\prime\prime\prime}(y,\varrho)|$ |′′$\leq|\ell^{\prime\prime}(y,\varrho)|,\quad\text{ for all }\varrho\in\mathbb{R}$ (3)
@@
$\widehat{\beta}^{(s)}$ $=\text{argmin}_{\beta\in\mathcal{B}}R[\mathbf{h}^{\intercal}\beta|Y,X,s]$
$\widehat{f}_{MR}^{(s)}$ $:=\text{argmin}_{f\in\mathcal{F}}R[f|Y,X,s;\widehat{w}_{s}]+\lambda^{\prime}_{s}\lVert f-\widehat{\beta}^{(s)\intercal}\mathbf{h}\rVert^{2}_{\mathcal{F}}$
$=\text{argmin}_{f\in\mathcal{F}:\lVert f-\widehat{\beta}^{(s)\intercal}\mathbf{h}\rVert\leq\nu^{\prime}}R[f|Y,X,s,\widehat{w}_{s}]$
@@
$\Delta_{MR,f,s}$ $:=2\eta L\bigg{(}\frac{2\nu^{\prime}\sqrt{\text{tr}(\mathbf{K}_{s})}}{n_{s}}+3\sqrt{\frac{\log(2/\delta)}{2n_{s}}}+$
$\frac{\lVert\widehat{\beta}^{(s)\intercal}\mathbf{h}(\mathbf{X}_{s})\rVert_{2}}{n_{s}}\bigg{)}$
where $\lVert\widehat{w}_{s}-w_{s}\rVert_{1}=\sum_{i\in N_{s}}|\widehat{w}_{s}(X_{i},Y_{i})-w_{s}(X_{i},Y_{i})|$ and $\mathbf{K}_{s}$ is the Gram matrix corresponding to the kernel function $k(\cdot,\cdot)$ associated with the space ${\cal F}$, $\mathbf{K}_{s,ij}=k(X_{i},X_{j}):i,j\in N_{s}$, and $\lVert\widehat{\beta}^{(s)\intercal}\mathbf{h}(\mathbf{X}_{s})\rVert_{2}=\sum_{i\in N_{s}}(\widehat{\beta}^{(s)\intercal}\mathbf{h}(X_{i}))^{2}$
@@
An alternative competitor one could suggest is to apply the DR estimator separately over every segment. Instead of employing a linear combination for the first stage, the DR learns an estimate $\widehat{f}^{(s)}_{p,\nu_{p,s}}\in{\cal F}$ trained on the unweighted empirical risk. Our comparison focuses on the functional complexity component of our bounds. They include a term analogous to our $\nu^{\prime}$ in their bound, which we denote as $\nu_{DR}^{\prime}$. However, they demonstrate $\nu_{DR}^{\prime}=\lVert f^{*(s)}_{\nu_{p,s}}-f^{*(s)}_{\nu_{q,s}}\rVert_{{\cal F}}+\mathcal{O}(n_{s}^{-1/4})$, where $f^{*(s)}_{\nu_{p,s}}$ represents the population optimizer of the unweighted problem. In contrast, our first stage estimator is expressed as $\nu^{\prime}=\lVert\beta^{*(s)\intercal}\mathbf{h}-f^{*(s)}_{\nu_{q,s}}\rVert_{\mathcal{F}}+\mathcal{O}(n_{s}^{-1/2})$, suggesting an improved rate for segments with small samples. Moreover if $\beta^{*(s)\intercal}\mathbf{h}$ proves to be a better approximation through similar segments to $f^{*(s)}_{\nu_{q,s}}$ then we can improve on the constant factor as well.
Remark 5.7 (Curse of dimensionality).
In this analysis, it’s important to acknowledge that our risk minimization is centered around comparing the estimator to the population minimizer within the function class of the Reproducing Kernel Hilbert Space (RKHS), thereby circumventing the curse of dimensionality. However, if one aims to minimize $R_{Q}[\cdot|s]$ across a broader function class we would expect to observe a dependence on the dimension of the features. Additionally, under covariate shift, the curse of dimensionality could potentially affect the estimation of the weights $\Delta_{MR,w_{s}}$. Nonetheless, our theoretical findings offer valuable insights into our method’s behavior.